# FactoFormer: Factorized Hyperspectral Transformers with Self-Supervised Pre-Training

### 11 Dec 2023 v2

@@ VI-B Effect of Factorization
In order to showcase the significance of factorized transformers and underscore their advantages in pre-training, we compare the performance with pre-training and without pre-training, i.e.,training from scratch in a fully supervised manner, for both FactoFormer and MAEST[14]. For this, we trained FactoFormer with a fully supervised setup by training the factorized transformers end-to-end from scratch for 300 epochs while searching for the best hyper-parameters. Accuracy of SpectralFormer[13] is considered as non-pre-trained MAEST[14] since the architecture of the classification encoder in MAEST[14] is the same as SpectralFormer[13]. Table IX shows the results evaluated on the Indian Pines dataset, and OA, AA and k are reported. FactoFormer outperforming MAEST[14] without pre-training shows that factorization facilitates learning more meaningful feature representations from both spectral and spatial dimensions. When we compare the improvement of accuracy when pre-trained, MAEST[14] shows a 2.39% improvement, whereas FactoFormer improves by 8.67%. This significant improvement shows that factorization enables more meaningful and effective pre-training. Moreover, pre-trained, FactoFormer outperforming pre-trained MAEST [14] by 7.15% proves the effectiveness of FactoFormer for hyperspectral data.
@@ VI-C Ablation Study
In this section, we conduct ablation studies to analyze the contribution of each component in FactoFormer. First, we investigate the advantage of using the factorized architecture over using them separately. Next, we examine the efficiency of using self-supervised pre-training compared to a fully supervised setting. Then we attempt pre-training with different masking ratios and different patch sizes. Finally, we analyze the computational efficiency of FactoFormer. Unless stated otherwise, we use the Indian Pines dataset for all our experiments.
Factorized Architecture: To showcase the effectiveness of the factorized transformers, we conducted separate pre-training and fine-tuning for each transformer. As presented in Table X, a significant drop in performance was observed when only one transformer was used. The factorized architecture allows the spectral and spatial transformers to learn feature representations through spatially and spectrally consistent masking strategies. When fine-tuning FactoFormer end-to-end, both transformers are initialized with separately pre-trained weights, which allows the model to benefit from the complementary features extracted from both transformers, leading to improved performance.
@@
Comparison to Masked Spatial-Spectral Transformer: To highlight the performance and the computational efficiently of FactoFormer we compare it with a masked spatial-spectral transformer. Initially we pre-train the model by masking spectral-spatial patches and reconstructing them as shown in Figure 7. We conducted the experiment by patching the input sample $S\times S\times B$ into non overlapping patches of size $1\times 1\times k$. Then the pre-trained encoder is used to fine-tune the network to perform HSI classification. We set $k$ to $10$ and conducted experiments using the Indian Pines dataset finding the best hyper-parameter setup. As shown in Table XI, FactoFormer performs significantly better with high accuracy and less than half the FLOPs when compared to the masked spatial-spectral transformer network.
Figure 7: An overview of masked spectral-spatial transformer which is pre-trained by masking random portion of spectral-spatial patches and reconstructing the masked regions.
@@
Masking Ratios: The masking ratio determines the number of visible tokens during pre-training, which is crucial for the model to learn and understand the patterns in the input data. However, setting this value too high or too low can negatively affect performance. In order to determine the optimal masking ratio for pre-training, we conducted a series of experiments using different masking ratios (0.5 to 0.8) for both spectral and spatial transformers. The results of this grid search are presented in Table XII. Based on this study, we selected a masking ratio of 0.7 for both transformers as the optimal value for fine-tuning the final model.
@@
Computational Efficiency: In order to compare the computational complexity of the proposed factorized pre-training approach with MAEST [14], we conducted a study using the same resources to compute the run times, and the average of five epochs was reported. Table XV shows the number of parameters and the pre-training run time per epoch for both FactoFormer and MAEST[14] methods. The autoencoder architecture of MAEST[14] requires heavy decoders to reconstruct the whole input, resulting in significantly high run-time and number of parameters. In contrast, we utilize a single linear layer prediction head to predict the original signals at the masked regions, substantially reducing the number of parameters. The sum of parameters in our spatial and spectral encoders is only one-third compared to MAEST[14]. The run times for both spatial and spectral transformers in our approach are approximately 7 seconds, even though the number of parameters is almost fourfold in the spatial Transformer. The reported times are similar due to the variation in the length of the input sequence. Compared to MAEST[14], the proposed FactoFormer reports a twelve times lower run-time, demonstrating that the factorized pre-training approach is computationally more efficient.
Impact of Dataset Size for Fine-Tuning: We also conducted a study to analyze the performance of the pretrained model by varying the size of training data used during fine-tuning. For this experiment, we used different percentages of data samples from the Indian Pines dataset, ranging from 20% to 100%. We compared FactoFormer with MAEST [14] and the results are shown in Figure 8. Notably, FactoFormer consistently outperforms MAEST [14] across all the different data percentages. As we increase the amount of data, the accuracy further improves. This analysis demonstrates that the model learns a better representation through the proposed factorized self-supervised pre-training.
@@
TABLE XV: Pre-training Computational Complexity Analysis
@@ VII Discussion
FactoFormer demonstrates state-of-the-art performance across all six datasets and outperforms other transformer based methods by a significant margin. The main advantage of FactoFormer is that it enables factorized self-attention and factorized self-supervised pre-training, resulting in improved performance compared to SpectralFormer [13] and MAEST [14]. We highlight the importance of utilizing both spectral and spatial features compared to focusing only on spectra with a more challenging downstream task like precise crop classification using the WHU-Hi dataset. We also show that our pre-training is computationally more efficient than MAEST [14] since we utilize a single linear layer prediction head as the decoder. Additionally, it’s worth noting that our models converge in fewer epochs during fine-tuning on all datasets, which further highlights the efficiency of the pre-training approach. Furthermore, our approach consistently outperforms MAEST [14] when fine-tuned with different percentages of training data.
Our pre-training approach shares the same limitation as other available techniques when dealing with data collected from various sensors. Hyperspectral imaging datasets exhibit variations in properties such as wavelength range, spectral resolution, the number of bands, and spatial resolution across various sensors. With our current approach in common with most other reported approaches, we need to pre-train the model every time we use a new dataset collected from a different sensor. For this reason, cross-domain hyperspectral image classification is drawing increasing attention [22], [23], [24]. In the case of self-supervised learning, pre-training in the language and vision domains has shown that pre-training on large datasets achieves state-of-the-art performance. Handling sensor variations in HSI and combining datasets for pre-training is an interesting direction for future research.
@@ VIII Conclusion
This paper introduced a novel factorized transformer architecture, FactoFormer, with self-supervised pre-training for hyperspectral data. The factorized architecture enables factorized self-attention, where attention is computed by individually focusing on spectral and spatial dimensions in each transformer, leading to the learning of salient spatial-spectral representations. The proposed masking strategy within the self-supervised pre-training pipeline has significantly contributed to better representation learning in the proposed model. FactoFormer achieves state-of-the-art performance while being computationally more efficient compared to the previous state-of-the-art transformer methods. In future research, we will address the problem of unsupervised domain adaptation by investigating how to transfer the learned representations between different hyperspectral datasets while utilizing common properties, which will enable pre-training on large hyperspectral datasets and fine-tuning on smaller datasets.

### 04 Jan 2024 v4

@@ VI-B Effect of Factorization
In order to showcase the significance of factorized transformers and underscore their advantages in pretraining, we compare the performance with pretraining and without pretraining, i.e.,training from scratch in a fully supervised manner, for both FactoFormer and MAEST[14]. For this, we trained FactoFormer with a fully supervised setup by training the factorized transformers end-to-end from scratch for 300 epochs while searching for the best hyper-parameters. Accuracy of SpectralFormer[13] is considered as non-pretrained MAEST[14] since the architecture of the classification encoder in MAEST[14] is the same as SpectralFormer[13]. Table IX shows the results evaluated on the Indian Pines dataset, and OA, AA and k are reported. FactoFormer outperforming MAEST[14] without pretraining shows that factorization facilitates learning more meaningful feature representations from both spectral and spatial dimensions. When we compare the improvement of accuracy when pretrained, MAEST[14] shows a 2.39% improvement, whereas FactoFormer improves by 8.67%. This significant improvement shows that factorization enables more meaningful and effective pretraining. Moreover, pretrained, FactoFormer outperforming pretrained MAEST [14] by 7.15% proves the effectiveness of FactoFormer for hyperspectral data.
@@ VI-C Ablation Study
In this section, we conduct ablation studies to analyze the contribution of each component in FactoFormer. First, we investigate the advantage of using the factorized architecture over using them separately. Next, we examine the efficiency of using self-supervised pretraining compared to a fully supervised setting. Then we attempt pretraining with different masking ratios and different patch sizes. Finally, we analyze the computational efficiency of FactoFormer. Unless stated otherwise, we use the Indian Pines dataset for all our experiments.
Factorized Architecture: To showcase the effectiveness of the factorized transformers, we conducted separate pretraining and fine-tuning for each transformer. As presented in Table X, a significant drop in performance was observed when only one transformer was used. The factorized architecture allows the spectral and spatial transformers to learn feature representations through spatially and spectrally consistent masking strategies. When fine-tuning FactoFormer end-to-end, both transformers are initialized with separately pretrained weights, which allows the model to benefit from the complementary features extracted from both transformers, leading to improved performance.
@@
Comparison to Masked Spatial-Spectral Transformer: To highlight the performance and the computational efficiently of FactoFormer we compare it with a masked spatial-spectral transformer. Initially we pretrain the model by masking spectral-spatial patches and reconstructing them as shown in Figure 7. We conducted the experiment by patching the input sample $S\times S\times B$ into non overlapping patches of size $1\times 1\times k$. Then the pretrained encoder is used to fine-tune the network to perform HSI classification. We set $k$ to $10$ and conducted experiments using the Indian Pines dataset finding the best hyper-parameter setup. As shown in Table XI, FactoFormer performs significantly better with high accuracy and less than half the FLOPs when compared to the masked spatial-spectral transformer network.
Figure 7: An overview of masked spectral-spatial transformer which is pretrained by masking random portion of spectral-spatial patches and reconstructing the masked regions.
@@
Masking Ratios: The masking ratio determines the number of visible tokens during pretraining, which is crucial for the model to learn and understand the patterns in the input data. However, setting this value too high or too low can negatively affect performance. In order to determine the optimal masking ratio for pretraining, we conducted a series of experiments using different masking ratios (0.5 to 0.8) for both spectral and spatial transformers. The results of this grid search are presented in Table XII. Based on this study, we selected a masking ratio of 0.7 for both transformers as the optimal value for fine-tuning the final model.
@@
Computational Efficiency: In order to compare the computational complexity of the proposed factorized pretraining approach with MAEST [14], we conducted a study using the same resources to compute the run times, and the average of five epochs was reported. Table XV shows the number of parameters and the pretraining run time per epoch for both FactoFormer and MAEST[14] methods. The autoencoder architecture of MAEST[14] requires heavy decoders to reconstruct the whole input, resulting in significantly high run-time and number of parameters. In contrast, we utilize a single linear layer prediction head to predict the original signals at the masked regions, substantially reducing the number of parameters. The sum of parameters in our spatial and spectral encoders is only one-third compared to MAEST[14]. The run times for both spatial and spectral transformers in our approach are approximately 7 seconds, even though the number of parameters is almost fourfold in the spatial Transformer. The reported times are similar due to the variation in the length of the input sequence. Compared to MAEST[14], the proposed FactoFormer reports a twelve times lower run-time, demonstrating that the factorized pretraining approach is computationally more efficient.
Impact of Dataset Size for Fine-Tuning: We also conducted a study to analyze the performance of the pretrained model by varying the size of training data used during fine-tuning. For this experiment, we used different percentages of data samples from the Indian Pines dataset, ranging from 20% to 100%. We compared FactoFormer with MAEST [14] and the results are shown in Figure 8. Notably, FactoFormer consistently outperforms MAEST [14] across all the different data percentages. As we increase the amount of data, the accuracy further improves. This analysis demonstrates that the model learns a better representation through the proposed factorized self-supervised pretraining.
@@
TABLE XV: Pretraining Computational Complexity Analysis
@@ VII Discussion
FactoFormer demonstrates state-of-the-art performance across all six datasets and outperforms other transformer based methods by a significant margin. The main advantage of FactoFormer is that it enables factorized self-attention and factorized self-supervised pretraining, resulting in improved performance compared to SpectralFormer [13] and MAEST [14]. We highlight the importance of utilizing both spectral and spatial features compared to focusing only on spectra with a more challenging downstream task like precise crop classification using the WHU-Hi dataset. We also show that our pretraining is computationally more efficient than MAEST [14] since we utilize a single linear layer prediction head as the decoder. Additionally, it’s worth noting that our models converge in fewer epochs during fine-tuning on all datasets, which further highlights the efficiency of the pretraining approach. Furthermore, our approach consistently outperforms MAEST [14] when fine-tuned with different percentages of training data.
Our pretraining approach shares the same limitation as other available techniques when dealing with data collected from various sensors. Hyperspectral imaging datasets exhibit variations in properties such as wavelength range, spectral resolution, the number of bands, and spatial resolution across various sensors. With our current approach in common with most other reported approaches, we need to pretrain the model every time we use a new dataset collected from a different sensor. For this reason, cross-domain hyperspectral image classification is drawing increasing attention [22], [23], [24]. In the case of self-supervised learning, pretraining in the language and vision domains has shown that pretraining on large datasets achieves state-of-the-art performance. Handling sensor variations in HSI and combining datasets for pretraining is an interesting direction for future research.
@@ VIII Conclusion
This paper introduced a novel factorized transformer architecture, FactoFormer, with self-supervised pretraining for hyperspectral data. The factorized architecture enables factorized self-attention, where attention is computed by individually focusing on spectral and spatial dimensions in each transformer, leading to the learning of salient spatial-spectral representations. The proposed masking strategy within the self-supervised pretraining pipeline has significantly contributed to better representation learning in the proposed model. FactoFormer achieves state-of-the-art performance while being computationally more efficient compared to the previous state-of-the-art transformer methods. In future research, we will address the problem of unsupervised domain adaptation by investigating how to transfer the learned representations between different hyperspectral datasets while utilizing common properties, which will enable pretraining on large hyperspectral datasets and fine-tuning on smaller datasets.